OpenStudy (anonymous):

3. 75.36= 3.14x^2(6) Simplify the equation and write it in standard form. Multiply each term in the equation by 100 to eliminate any decimals, if necessary.

OpenStudy (anonymous):

the 3 is just the # of questain not part of it

OpenStudy (turingtest):

this question is still illegible...

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

what is that 6 doint there?

OpenStudy (anonymous):

well im doing an assingment and hold on its easyer to just post the directions i had i prolly did it wrong

OpenStudy (anonymous):

Measure and record the diameter and height of the cylindrical object you have chosen in inches. Round to the nearest whole number. Apply the formula of a right circular cylinder (V = r2h) to find the volume of the object. (Note: Be sure to find the radius from the diameter measurement by dividing by 2.) Now suppose you knew the volume of this object and the relation of the height to the radius, but did not know the radius. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the radius. Rewrite the formula using the variable x for the radius. Substitute the value of the volume found in step 2 for V and express the height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the radius, then the expression for the height will be (x + 4). ... and iv done this 1. D= 4in. H= 6in. 2. 3.14 (2^2)6 / 3.14(4)(6)= 75.36 3. 75.36= 3.14x^2(6) 4.

OpenStudy (anonymous):

\[75.36= 3.14x^2\] \[7536=314x^2\]

OpenStudy (anonymous):

am i doing it right?